A charged oil-drop of radius 1.3 x 10-4cm is prevented from falling under gravity by the vertical field between two horizontal plates charged to a difference of potential of 8340 volts. The distance between the plates is 1.6 cm, and the density of oil is 920kgm-3. Calculate the magnitude of the charge on the drop.
1
Expert's answer
2020-11-06T10:18:12-0500

Volume of oil:

"V=\frac{4}{3}\times \pi \times R^3"

"V=\frac{4}{3}\times 3.14 \times (1.3\times10^{-6})^3=\frac{4}{3}\times 3.14 \times 2.197\times10^{-18}"

"V=9.2\times 10^{-18}m^3"

Mass of oil:

"\rho=\frac{m}{V}"

"m=\rho \times V=920\times 9.2 \times 10^{-18}=8.464 \times 10^{-15}"

Gravity force for oil:

"F_{g}=m\times g = 8.464 \times 10^ {-15}\times9.8 = 8.3 \times10^{-14}"

Electric field between plates:

"E=\frac{U}{d}=\frac{8340}{0.016}=521250"

Charge of oil:

"q=\frac{F}{E}=\frac{8.3\times10^{-14}}{521250}=1.6\times10^{-19}"